Determine the given and compute the test statistic of the problem below using Central Limit Theorem, and construct the rejection region for each.


A certain group of welfare recipients receives relief goods with a mean amount of Php 500 per week. A random sample of 75 recipients is surveyed and found that the mean amount of relief goods they received in a week is Php 600 and a standard deviation of Php50.00 . Test the claim at 1% level of significance is not Php 500 per week and assume that the population is approximately normally distributed.


1
Expert's answer
2022-05-10T22:57:45-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=500"

"H_a:\mu\not=500"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=74" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.643913."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.643913\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{600-500}{50/\sqrt{75}}\approx17.3205"


5. Since it is observed that "|t| =17.3205> 2.643913=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=74" degrees of freedom, "t=17.3205" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 500, at the "\alpha = 0.01" significance level.